OpenStudy (logan13):

Barkley has a box shaped like a cylinder that he will use to package a ball for a gift. He wants to fill the remainder of the box with packing foam. Which formulas will he need to use to find the volume of packing foam he will need? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.)

OpenStudy (logan13):

http://prntscr.com/adxkmr

OpenStudy (nubeer):

i think it would be A

OpenStudy (anonymous):

A, because a circle's area is...\[\pi r ^{2}\] If you were to stretch that circle up into a cylinder you are basically multiplying by the height hence....\[\pi r ^{2} h\]

OpenStudy (anonymous):

Also don't forget since it's 3 dimensional, it will be measured in cubic units.

OpenStudy (logan13):

its multiple choice

OpenStudy (anonymous):

I know, but for future reference :)

OpenStudy (logan13):

so what other answer is it

OpenStudy (anonymous):

A :)

OpenStudy (logan13):

A and

OpenStudy (anonymous):

Oh sorry and for the spherical ball that would be...\[\frac{ 4 }{ 3 }\pi r ^{3}\]

OpenStudy (anonymous):

so A and C. My apologies I didn't read the question properly.